OpenStudy (anonymous):

There were 150 people at the Junior-Senior dance Sasha and her friends attended. Junior tickets were $2.00 each and Senior tickets were $3.50 each. The total receipts for the dance were $405. How many Juniors bought tickets? 70 80 90 100

OpenStudy (anonymous):

pleasee heelllp. ugh.

OpenStudy (radar):

Let j = the number of juniors Let s = the number of seniors Now from the information given we can develop the following two equations. 1. j + s = 150 Now for the money. 2, You see we now have a system of equations, 2 equatios, 2 unknownhs, j($2.00) + s(#3.50)=$405

OpenStudy (radar):

There are two common methods to solve this system of linear equations. Which method do you prefer?

OpenStudy (anonymous):

assume all seniors bought tickets @ $3.50. So $3.50 * 150 = 525.

OpenStudy (anonymous):

then subtract the given total from 525

OpenStudy (anonymous):

you should have $120 difference, now the difference in a junior ticket is $1.50. so how many $1.50 can you get from $120

OpenStudy (anonymous):

your answer is 80

OpenStudy (anonymous):

I am not sure what method to use on this..

OpenStudy (radar):

The methods usually used are "substitution" or "elimination"

OpenStudy (anonymous):

I used the elimination method

OpenStudy (radar):

Lets use substitution: From equation 1. s=150-j we now substitute the s value in equation 2. j($2.00) + (150-j)($3.50)=$405 2j+(3.5)(150)-3.5j=$405 -1.5j=-120 j=80

OpenStudy (radar):

As you can see either method the results are the same.